OpenStudy (anonymous):

5+ (square root) 3 divided by 2-(square root) 3

OpenStudy (anonymous):

\[=(5+\sqrt{3})*(2+\sqrt{3})/(2-\sqrt{3})*(2+\sqrt{3})=13+7\sqrt{3}\]

OpenStudy (anonymous):

thank you